### v5

@@ 2.1 Natural Selection and Embodiment
We assume a naturally selected, embodied organism. Every aspect of the environment is a set of facts [^5] Some may object, pointing out that this ignores composition. However, the application of a function is a fact regardless of whether it takes another function as input., hence the body of an organism that persists across multiple states must be a set of programs. We call this set a vocabulary, and it is formalised in definition 2 (see appendix [40]).
Piccinini [34, 35] divides computation into abstract and concrete sorts. Abstract computations mean or represent whatever we say they do. In contrast concrete computations are the actions of physical systems. The declarative programs we describe are concrete. Each captures a single point of difference between physical states. As such no two organisms have the same vocabulary because that would mean they are the same body.
@@
This might seem inconvenient, as intuitively we might prefer to think about bodies as an abstract ‘type’. However, doing so would undermine any claim we might make, as earlier results in artificial intelligence (AI) show [39, 43]. The meaning of abstractions depends upon interpretation, much like the behaviour of software is determined by the abstraction layer that “interprets” it. At its least abstract software is machine code [^6] A “word” of machine code triggers a mechanistic process hardwired in the CPU by the human design. For example a word may copy the 32 bit value stored at memory address $X$ into the 32 bit register $Y$, which the next line of machine code “adds” to register $Z$ by looping over each bit in an “adder” circuit.. Machine code is interpreted by hardware which determines every aspect of what that machine code does. A word of machine code is a mechanical trigger that only “means” whatever we have designed the hardware to do when we input that word. That hardware is an abstraction layer in which the software exists (including “higher level” abstraction layers like the Python interpreter). In other words, what we call software is nothing more than the state of hardware [36]. This has undermined all but the most subjective of claims regarding the behaviour of theorised, software superintelligence [44, 45, 46]. It is a flaw in the very idea of intelligent software. The distinction between software mind and hardware embodiment is subsequently called computational dualism [36], because it is reminiscent of how traditional Cartesianism conceived of a mental substance distinct from physical.
@@
If we are to dissolve the hard problem by starting with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states, then we cannot presuppose organisms use particular abstractions. Instead we must explain why particular abstractions are formed, in terms of contentless states, as the product of interaction between an organism and its environment [33, 26, 47]. This process is sometimes called “relevance realisation” [48, 49, 50].
@@
Some have argued enactivism is incompatible with computation because the set of possible abstractions that must be searched for relevance realisation is intractable [51]. For tractability, one must first isolate a “small world” of relevant information from the “big world” of all information [52]. Part of the reason we can address this is because we formalise embodiment, in concrete terms from the level of irreducible physical states. Embodiment implies finite resource constraints, so we adopt one more axiom from [41].
Axiom 3: All aspects of the environment are spatially extended [41].
@@
This means that a body is realised by only a finite number of states (see def. 2). A body is a vocabulary $\mathfrak{v}\subset P$ with finitely many elements. $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}$ of interaction between body and environment. A statement in this language is an aspect of the environment. A set of “programs” from our vocabulary. Statements have truth conditions with respect to environmental states, and every statement has an extension. The extension of a statement is the set of all statements which are supersets of the first statement (the set of all other statements by which the first statement is implied). Extension is important we can relate statements by their truth conditions, forming a lattice. Most importantly, we avoid the distinction between software and hardware and so avoid computational dualism.
@@
Embodiment is dispositional. As natural selection evolves a vocabulary, it begins to isolate a “small world” relevant to survival [51, 41, 53], from the intractable “big world” $P$ of all information.
@@ 2.2 Self-Organising Systems as Self and World Constraints
Using our embodied formal language $L_{\mathfrak{v}}$, we can talk about computation with inputs and outputs by treating everything as embodied statements embedded and enacted within the environment. Given an input $i\in L_{\mathfrak{v}}$, the set of all possible outputs is the extension $E_{i}$ of that input. This is because if $i$ is realised by the environment, then the environment is constrained to those states that realise $i$, which constrains what other statements are realised. If behaviour is “motivated”, then statements have valence determined by natural selection and only a subset of the statements a body could make are “fit”. Hence an organism will self-organise to express some statements, but not others. Though environmental states are contentless and their only significance is in how they relate to one another, natural selection must make them attractive or repulsive to a self-organising system. This is formalised by the $\mathfrak{v}$-task as in definition 3777A $\mathfrak{v}$-task is called because a $\mathfrak{v}$-task because it exists in the context of a vocabulary $\mathfrak{v}$.. If $\mathfrak{v}$ is the vocabulary of a body, and $\mathfrak{v}$-task $\mu=\langle I_{\mu},O_{\mu}\rangle$ is fit behaviour, then $I_{\mu}$ is all the statements that body can express in which it is possible to remain fit, and $O_{\mu}$ is all the statements that body can express in which it remains fit. The extension $E_{I_{\mu}}$ of $I_{\mu}$ would be every output that it is possible to choose given the inputs $I_{\mu}$, but only a subset of those $O_{\mu}\subset E_{I_{\mu}}$ are correct outputs.
@@ 2.4 Learning
If an organism remains alive, then its history is an ostensive definition of “fit” self-organising behaviour. That history is expressed as a $\mathfrak{v}$-task $\mathfrak{h}$, where each input and output in $\mathfrak{h}$ is an interaction between organism and environment. If the set of all fit behaviour is a task $\mu$, then the organism remains fit then $\mathfrak{h}\sqsubset\mu$.
“I survive therefore my model is viable.” - Mark Solms [54]
@@
“The best model of the world is the world itself” - Rodney Brooks [55]
Earlier work [38] showed that the optimal strategy to adapt as fast as possible [^9] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour. is to prefer “weaker” policies, meaning those with larger extensions. The policy which generated outputs $O$ given inputs $I$ is most efficiently identified by constructing a policy $\pi$ such that $\pi$ generates $O$ from $I$, and $\pi$ implies the weakest constraint that can be implied while still generating $O$ from $I$. This means a system that does so will construct fit policies from a shorter history [39] than one that does not (see def. 5 and proofs in appendix [40]) [^10] The effect of insufficiently weak representations can be observed in large language models like GPT-4 [56, 57], where they fail to make consistent statements about an object because represent it as multiple objects..
@@ 3 Relevance Realisation Through Causal Learning
All organisms (see definition 6) must have preferences, because they will make some decisions and not others. Behaviour implies policies, and every policy implies tasks, so we can define preferences as an binary relation over tasks.
@@
Natural selection favours adaptability, hence it prefers organisms that optimise for weak policies, which we call weak policy optimisation (w-maxing). By constructing fit policies, an organism divides the world into relevant objects and properties. A weaker policy implies all more specific versions of itself, meaning those that more tightly constrain outputs by having a smaller extension. Hence policy learning implies a lattice of policies that vary in weakness. Every fit policy would be a “causal identity” for something relevant, for example a specific object like “this coffee” or a weaker and more general concept like “all coffee”.
@@
Organisms interact when they “affect” one another (see definition 9) in the causal or physical rather than psychological sense of the word. Now, Artificial Intelligence (AI) and machine learning [58, 59, 60, 61] are concerned with engineering adaptive agents. In that context, causality has become a mainstream topic of research. Causal learning is demonstrably necessary to thrive in an interactive setting [62, 63]. Where the causal graph is known in advance (for example if measuring the efficacy of medical interventions), this issue could be resolved using causal language [64] to represent an intervention from outside the system the graph describes.
@@
To illustrate [39], suppose Bob is attempting to learn and predict the environment. Bob has observed Alice wearing a raincoat only when it rains, and Bob has seen rain only when he observed Alice wearing a raincoat. If we represent the raincoat observation with a binary variable $O\in\{true,false\}$ and the advent of rain with a variable $R\in\{true,false\}$, and if Bob is a Bayesian, then Bob’s observations will lead to the conclusion that $p(R=true\mid O=true)=1$. This means Bob believes that if Alice wears a raincoat, then it must be raining. Now lets permit Bob to interact with its environment. Assume Bob wants it to rain. As $p(R=true\mid O=true)=1$, Bob may conclude that forcing $O=true$ by holding a gun to Alice’s head and demanding Alice wear a raincoat will cause it to rain. This is absurd. $O=true$ does not represent the event $q=$“I coerced Alice into wearing a raincoat”, but an entirely different event $v=$“Alice decided to put on a raincoat for the same reason I have observed Alice wearing a raincoat in the past”. To accurately represent the environment, we need to represent that $p(R=true\mid q)=p(R=true)\neq p(R=true\mid v)=1$, meaning both $q$ and $v$. To illustrate visually, we started with the acyclic graph
$O$$R$
and our intervention disconnected rain from the choice of clothing:
$O$$R$
This can be resolved by introducing a “do” operator that we apply to a variable $O$ to obtain $do[O=o]$, to represent the fact that an agency from outside the system has intervened to assign a value to $O$, so that we can represent $p(R=true\mid do[O=true])=p(R=true)\neq p(R=true\mid O=true)=1$. Thus, the aforementioned $q$ is equivalent to $do[O=true]$, while $v$ is equivalent to $O=true$.
3.0.1 The Psychophysical Principle of Causality
@@
There are two ways to learn causal graphs. One can either assume a set of variables and learn relations between them, or assume relations and learn the objects that fit them [39]. We do the latter. Rather than assuming objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed according to what causes valence. We call this the psychophysical principle of causality. This is what will take us from attraction to and repulsion from contentless physical states, through lower order thoughts to higher order meta representations. We start with the complete absence of content [65], and once a system learns policies there are distinct “contents”. Policies serve to classify objects and properties. Death grounds meaning.
To develop this idea, we first show adding additional nodes to a causal graph is an adequate substitute for the “do” operator [66]. Assume we again have Bob who constructs a causal graph of the environment. Assume Alice exists in that environment. From Bob’s perspective, Alice is just a part of the environment represented by a variable $A$ in Bob’s causal graph.
$O$$A$$R$
Now assume Bob observes Alice taking an action that changes an aspect of the environment, represented by the variable $O$ (for example, Bob observes Alice putting on the raincoat). From Bob’s perspective, Alice’s action assigns values to variables $A$ and $O$. There is no need to involve a do operator in this scenario because we can already represent that $p(R=true\mid A=x,O=true)=p(R=true)\neq p(R=true\mid A=y,O=true)=1$ (because Alice is part of the causal graph). Likewise, we can add variables for interventions of Bob [39].
$O$$B$$A$$R$
@@ 4 Relevant Causal Identities
The optimal choice of policy for adaptability is the weakest because it is implied by more statements in an embodied language. That language is dispositional [67], shaped by natural selection, and so a weaker policy is likely to be a fit policy. w-maxing will correctly identify cause and effect where relevant, simplifying the entire environment into objects and properties relevant to the organism’s motivations.
Assume a vocabulary $\mathfrak{v}_{\mathfrak{b}}$ belonging to an organism $\mathfrak{b}$ (Bob). A “cause” in the context of this formalisation is not a variable but a statement $l\in L_{\mathfrak{v}_{\mathfrak{b}}}$. The raincoat example would involve $obs,rain\in L_{\mathfrak{v}_{\mathfrak{b}}}$ such that:
[…]
$int$$obs$$rain$
So long as $obs\neq int$, the intervention can be differentiated from the passive observation (see definition 10).
This being the case, any set $c\subseteq int-obs$ could be used to identify the party undertaking the intervention, which is why $c$ is referred to as a “causal identity”. It distinguishes the intervention $int$ from the passively observed effect $obs$, like reafference in living organisms. However, the above only considers one intervention. A weaker or more general causal identity would be one that is shared by more interventions.
For example, suppose the inputs Alice has been subject to are $I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}$. These can be divided into those in which Bob affected Alice $I^{\mathfrak{b}}_{\mathfrak{a}}$ and those in which Bob did not $I^{\lnot\mathfrak{b}}_{\mathfrak{a}}=I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}-I^{\mathfrak{b}}_{\mathfrak{a}}$. Alice can construct a causal identity $b$ for Bob corresponding to interventions $INT=I^{\mathfrak{b}}_{\mathfrak{a}}$ and observations $OBS=I^{\lnot\mathfrak{b}}_{\mathfrak{a}}$ (see definition 11).

### v6

@@ 2.1 Natural Selection and Embodiment
We assume a naturally selected, embodied organism. Every aspect of the environment is a set of facts [^9] Some may object, pointing out that this ignores composition. However, the application of a function is a fact regardless of whether it takes another function as input., hence the body of an organism that persists across multiple states must be a set of programs. We call this set a vocabulary (formally defined in the appendix Bennett (2025a) as well as the summary at the beginning of this chapter).
@@
The meaning of abstractions depends upon interpretation, much like the behaviour of software is determined by the abstraction layer that “interprets” it. At its least abstract software is machine code [^10] A “word” of machine code triggers a mechanistic process hardwired in the CPU by the human design. For example a word may copy the 32 bit value stored at memory address $X$ into the 32 bit register $Y$, which the next line of machine code “adds” to register $Z$ by looping over each bit in an “adder” circuit.. Machine code is interpreted by hardware which determines every aspect of what that machine code does. A word of machine code is a mechanical trigger that only “means” whatever we have designed the hardware to do when we input that word. That hardware is an abstraction layer in which the software exists (including “higher level” abstraction layers like the Python interpreter). As stated earlier, what we call software is nothing more than the state of hardware Bennett (2024b). This has undermined all but the most subjective of claims regarding the behaviour of theorised, software superintelligence Hutter (2010); Leike and Hutter (2015, 2018). It is a flaw in the very idea of intelligent software. The distinction between software mind and hardware embodiment is subsequently called computational dualism Bennett (2024b), because it is reminiscent of how traditional Cartesianism conceived of a mental substance distinct from physical.
@@
If we start with the embodied biological organism instead of the abstract mental states, then we cannot presuppose organisms use particular abstractions. Instead we must explain why particular abstractions are formed, in terms of contentless states, as the product of interaction between an organism and its environment Thompson (2007); Varela et al. (2016); Rolla and Figueiredo (2021). This process is sometimes called “relevance realisation” Vervaeke et al. (2012); Vervaeke and Ferraro (2013b, a), and this shift in perspective is key for our argument here. Indeed, the idea that conscious experiences depend upon interactions with the environment has been defended by a longstanding tradition in embodied cognition and enactivism Thompson (2007).
@@
For example, some have argued enactivism is incompatible with computation at a fundamental physical level, because the set of possible abstractions that must be searched for relevance realisation is intractable Jaeger et al. (2024). For tractability, one must first isolate a “small world” of relevant information from the “big world” of all information Savage (1954). To address this we formalise embodiment, in concrete terms from the level of irreducible physical states. Embodiment implies finite resource constraints Bennett (2022b, 2025a); Ciaunica (2025), hence we adopt one more axiom.
Axiom 3: All aspects of the environment are spatially extended Bennett (2024c).
@@
This means that a body is realised by only a finite number of states. See the ESM for the self contained definition, and the Stack Theory appendix Bennett (2025a) for the full development. A body is a vocabulary $\mathfrak{v}\subset\mathcal{P}$ with finitely many elements. $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}$ of interaction between body and environment. A statement in this language is an aspect of the environment. A set of “programs” from our vocabulary. Statements have truth conditions with respect to environmental states, and every statement has an extension. The extension of a statement is the set of all statements which are supersets of the first statement (the set of all other statements by which the first statement is implied). Extension is important we can relate statements by their truth conditions, forming a lattice. Most importantly, we avoid the distinction between software and hardware and so avoid computational dualism.
@@
To summarise. Embodiment is dispositional Bennett (2024b, 2025b). As natural selection evolves a vocabulary, it begins to isolate a “small world” relevant to survival Jaeger et al. (2024); Bennett (2024c); Friston et al. (2023), from the intractable “big world” $\mathcal{P}$ of all information.
@@ 2.2 Self-Organising Systems as Self and World Constraints
Using our embodied formal language $L_{\mathfrak{v}}$, we can talk about computation with inputs and outputs by treating everything as embodied statements embedded and enacted within the environment. Given an input $i\in L_{\mathfrak{v}}$, the set of all possible outputs is the extension $E_{i}$ of that input. This is because if $i$ is realised by the environment, then the environment is constrained to those states that realise $i$, which constrains what other statements are realised. If behaviour is “motivated”, then statements have valence determined by natural selection and only a subset of the statements a body could make are “fit”. Hence an organism will self-organise to express some statements, but not others. Crucially, though environmental states are contentless they imply content by how they relate to one another. As states transition from one to another this preserves some aspects of the environment, and deletes others (which would happen when an aspect is made of facts that don’t hold at a particular moment in time, given the state at that time). As argued elsewhere, the environment preserves that which preserves itself Bennett (2025a). It follows that natural selection must make states attractive or repulsive to a self-organising system if that self-organising system is to be deemed ‘fit’. This is formalised by the $\mathfrak{v}$-task as in definition 3111111A $\mathfrak{v}$-task is called a $\mathfrak{v}$-task because it is defined relative to a vocabulary $\mathfrak{v}$.. If $\mathfrak{v}$ is the vocabulary of a body, and $\mathfrak{v}$-task $\mu=\langle I_{\mu},O_{\mu}\rangle$ is fit behaviour, then $I_{\mu}$ is all the statements that body can express in which it is possible to remain fit, and $O_{\mu}$ is all the statements that body can express in which it remains fit. The extension $E_{I_{\mu}}$ of $I_{\mu}$ would be every output that it is possible to choose given the inputs $I_{\mu}$, but only a subset of those $O_{\mu}\subset E_{I_{\mu}}$ are correct outputs.
@@ 2.4 Learning
If an organism remains alive, then its history is an ostensive definition of “fit” self-organising behaviour. That history is expressed as a $\mathfrak{v}$-task $\mathfrak{h}$, where each input and output in $\mathfrak{h}$ is an interaction between organism and environment. If the set of all fit behaviour is a task $\mu$, and the organism remains fit throughout its history, then $\mathfrak{h}\sqsubset\mu$.
“I survive therefore my model is viable.” - Mark Solms Solms (2021)
@@
“The best model of the world is the world itself” - Rodney Brooks Dreyfus (2007)
Earlier work Bennett (2023c) showed that the optimal strategy to adapt as fast as possible [^13] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour. is to prefer “weaker” policies, meaning those with larger extensions. The policy which generated outputs $O$ given inputs $I$ is most efficiently identified by constructing a policy $\pi$ such that $\pi$ generates $O$ from $I$, and $\pi$ implies the weakest constraint that can be implied while still generating $O$ from $I$. This means a system that does so will construct fit policies from a shorter history Bennett (2023a) than one that does not. Full proofs are in the ESM. [^14] The effect of insufficiently weak interpretations can be observed in large language models like GPT-3 and GPT-4 Floridi and Chiriatti (2020); Bennett (2023b), where they can fail to make consistent statements about an object because they represent it as multiple objects.
@@ 3 Relevance Realisation Through Causal Learning
All organisms must have preferences, meaning some futures are better for them than others. Formally, we model an organism as an embodied learner defined by an embodied vocabulary, a viability task, and a preference ordering over tasks. The full tuple definition is in the ESM. Behaviour implies policies, and every policy implies tasks, so preferences can be formalised as a binary relation over tasks.
@@
Natural selection favours adaptability, hence it prefers organisms that optimise for weak correct policies, which is referred to elsewhere as weakness maximisation (w-maxing) Bennett (2025c). By constructing fit policies, an organism divides the world into relevant objects and properties. A weaker policy implies all more specific versions of itself, meaning those that more tightly constrain outputs by having a smaller extension. Hence policy learning implies a lattice of policies that vary in weakness. Every fit policy would be a “causal identity” for something relevant, for example a specific object like “this coffee” or a weaker and more general concept like “all coffee”.
@@
Organisms interact when they can affect one another in the causal or physical rather than psychological sense of the word. Here affect means one organism’s interventions can change the other organism’s inputs, hence changing which outputs are viable. The formal definition we use is in the ESM. Now, AI and machine learning Russell and Norvig (2020); Bishop (2006); Sutton and Barto (2018); Goertzel (2014) are concerned with engineering adaptive agents. In that context, causality has become a mainstream topic of research. Causal learning is demonstrably necessary to thrive in an interactive setting Pearl and Mackenzie (2018); Richens and Everitt (2024). Where the causal graph is known in advance (for example if measuring the efficacy of medical interventions), this issue could be resolved using causal language Pearl (2009) to represent an intervention from outside the system the graph describes.
@@
To illustrate Bennett (2023a), suppose Bob is attempting to learn and predict the environment. Bob has observed Alice wearing a raincoat only when it rains, and Bob has seen rain only when he observed Alice wearing a raincoat. If we represent the raincoat observation with a binary variable $O\in\{true,false\}$ and the advent of rain with a variable $R\in\{true,false\}$, and if Bob is a Bayesian, then Bob’s observations will lead to the conclusion that $p(R=true\mid O=true)=1$. This means Bob believes that if Alice wears a raincoat, then it must be raining. Now let us permit Bob to interact with his environment. Assume Bob wants it to rain. As $p(R=true\mid O=true)=1$, Bob may conclude that forcing $O=true$ by holding a gun to Alice’s head and demanding Alice wear a raincoat will cause it to rain. This is absurd. $O=true$ does not represent the event q=q=“I coerced Alice into wearing a raincoat”, but an entirely different event v=v=“Alice decided to put on a raincoat for the same reason I have observed Alice wearing a raincoat in the past”. To accurately represent the environment, we need to represent that $p(R=true\mid q)=p(R=true)\neq p(R=true\mid v)=1$, meaning both $q$ and $v$. To illustrate visually, we started with the acyclic graph OORR and our intervention disconnected rain from the choice of clothing: OORR This can be resolved by introducing a “do” operator that we apply to a variable $O$ to obtain $do[O=o]$, to represent the fact that an agency from outside the system has intervened to assign a value to $O$, so that we can represent $p(R=true\mid do[O=true])=p(R=true)\neq p(R=true\mid O=true)=1$. Thus, the aforementioned $q$ is equivalent to $do[O=true]$, while $v$ is equivalent to $O=true$.
3.1 The Psychophysical Principle of Causality
@@
There are two ways to learn causal graphs. One can either assume a set of variables and learn relations between them, or assume relations and learn the objects that fit them Bennett (2023a). We do the latter. Rather than assuming objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed according to what causes valence. We call this the psychophysical principle of causality. This is what will take us from attraction to and repulsion from contentless physical states, through lower order thoughts to higher order meta representations. We start with the complete absence of content Thompson (2016), and once a system learns policies there are distinct “contents”. Policies serve to classify objects and properties. Death grounds meaning.
To develop this idea, we first show adding additional nodes to a causal graph is an adequate substitute for the “do” operator Dawid (2002). Assume we again have Bob who constructs a causal graph of the environment. Assume Alice exists in that environment. From Bob’s perspective, Alice is just a part of the environment represented by a variable $A$ in Bob’s causal graph. OOAARR Now assume Bob observes Alice taking an action that changes an aspect of the environment, represented by the variable $O$ (for example, Bob observes Alice putting on the raincoat). From Bob’s perspective, Alice’s action assigns values to variables $A$ and $O$. There is no need to involve a do operator in this scenario because we can already represent that $p(R=true\mid A=x,O=true)=p(R=true)\neq p(R=true\mid A=y,O=true)=1$ (because Alice is part of the causal graph). Likewise, we can add variables for interventions of Bob Bennett (2023a). OOBBAARR
@@ 4 Relevant Causal Identities
The optimal choice of policy for adaptability is the weakest because it is implied by more statements in an embodied language. That language is dispositional Bennett and Maruyama (2022), shaped by natural selection, and so a weaker policy is likely to be a fit policy. w-maxing will correctly identify cause and effect where relevant, simplifying the entire environment into objects and properties relevant to the organism’s motivations.
Assume a vocabulary $\mathfrak{v}_{\mathfrak{b}}$ belonging to an organism $\mathfrak{b}$ (Bob). A “cause” in the context of this formalisation is not a variable but a statement $l\in L_{\mathfrak{v}_{\mathfrak{b}}}$. The raincoat example would involve $obs,rain\in L_{\mathfrak{v}_{\mathfrak{b}}}$ such that: obs ↔“Alice put on a raincoat” and rain ↔“It rained” obsobs and rainrain have truth values in accord with the definition. As in the earlier example Bob’s passive observation implies $p(rain\mid obs)=1$. The statement $obs=\textit{``Alice put on a raincoat''}$ can be made true by either passive observation or intervention. However, the statement which is true in the case of intervention not only obsobs, but $int\in L$ such that $obs\subseteq int$ and: int ↔“Alice is wearing a raincoat because of Bob’s actions” $\mathrm{int}$$\mathrm{obs}$$\mathrm{rain}$ So long as $obs\neq int$, the intervention can be differentiated from the passive observation at the more abstract representational scale of the vocabulary Bennett (2024a).
This being the case, any set $c\subseteq int-obs$ could be used to identify the party undertaking the intervention, which is why $c$ is referred to as a “causal identity”. It distinguishes the intervention intint from the passively observed effect obsobs, like reafference in living organisms. However, the above only considers one intervention. A weaker or more general causal identity would be one that is shared by more interventions.
A formal proposition shows that, when intervention episodes share at least one distinguishing signature, a w-max learner selects an inclusion minimal causal identity. In plain English, it learns the least committal intervention tag that still separates intervention from matched observation. See Box 3 for the sketch and see the ESM propositions on causal identity minimality for the full proof.
If a statement $l$ holds in both observation and intervention episodes, then $l$ alone cannot distinguish them. If $c\in\mathcal{C}(INT,OBS)$, then $l\cup c$ is true on every $int\in INT$ and false on every $obs\in OBS$. So $l\cup c$ plays the role of a do tag at the abstraction scale of $L_{\mathfrak{v}_{\mathfrak{a}}}$.
For example, suppose the inputs Alice has been subject to are $I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}$. These can be divided into those in which Bob affected Alice $I^{\mathfrak{b}}_{\mathfrak{a}}$ and those in which Bob did not $I^{\lnot\mathfrak{b}}_{\mathfrak{a}}=I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}-I^{\mathfrak{b}}_{\mathfrak{a}}$. Alice can construct a causal identity $b$ for Bob corresponding to interventions $INT=I^{\mathfrak{b}}_{\mathfrak{a}}$ and observations $OBS=I^{\lnot\mathfrak{b}}_{\mathfrak{a}}$. The full formalism for building these sets from a history is in the ESM.